# An Experimental Evaluation of Imputation Models for Spatial-Temporal Traffic Data

### v1

@@ Abstract
Traffic data imputation is a critical preprocessing step in intelligent transportation systems, enabling advanced transportation services. Despite significant advancements in this field, selecting the most suitable model for practical applications remains challenging due to three key issues: 1) incomprehensive consideration of missing patterns that describe how data loss along spatial and temporal dimensions, 2) the lack of test on standardized datasets, and 3) insufficient evaluations. To this end, we first propose practice-oriented taxonomies for missing patterns and imputation models, systematically identifying all possible forms of real-world traffic data loss and analyzing the characteristics of existing models. Furthermore, we introduce a unified benchmarking pipeline to comprehensively evaluate 10 representative models across various missing patterns and rates. This work aims to provide a holistic understanding of traffic data imputation research and serve as a practical guideline.
Index Terms:
traffic data imputation, spatial-temporal graph series, experimental evaluation.
@@ I Introduction
Traffic data is often represented with a graph of spatial-temporal series, where each node in the graph represents a traffic sensor continuously collecting observation series and edges in the graph describe the relationships between sensors. Traffic data holds a prominent role in Intelligent Transportation Systems (ITS), supporting key services in ITS, such as real-time traffic display, traffic data prediction, and traffic signal control. Despite the increasing deployment of sensors to collect real-time traffic information, missing data remains a persistent issue due to various factors like equipment malfunctions, power outages, and network failures. These missing data hinder ITS to accurately perceive traffic changes over time, resulting in the performance decline of transportation services. Therefore, traffic data imputation is a crucial preprocessing step in ITS, ensuring the quality of traffic data.
@@
Numerous models have been proposed to solve this problem, which primarily focus on addressing two challenges: 1) Effectively capturing the dynamics of traffic data with missing values across both spatial and temporal dimensions. 2) Training models with the absence of ground truth for the missing data. Beyond specialized traffic imputation models, general multivariate time series imputation models can also be adapted for this problem. Despite significant progress in this field, choosing an optimal model for specific real-world scenarios in ITS remains challenging due to three factors:
    Missing patterns are not comprehensive. In real-world scenarios, missing patterns in traffic data, which reflect how data loss along the spatial and temporal dimensions, exhibit significant diversity. For example, network failures may cause random traffic data loss, sensor malfunctions can result in continuous traffic data loss in the temporal dimension, and power outages may lead to sustained traffic data loss across both spatial and temporal dimensions. However, most existing studies consider only a limited subset of these patterns, leaving the performance of their models on real-world traffic data with other missing patterns largely unexplored.
    Datasets are not standardized. Existing studies often test their imputation models on different datasets, with model designs tailored to the specific characteristics of each datasets. There is a significant lack of empirical studies comparing different models on a standardized set of datasets, especially for spatial-temporal traffic data. Additionally, while both traffic flow and traffic speed are critical for ITS, few studies have conducted comprehensive experiments on both simultaneously.
@@
    Evaluations are not sufficient. Firstly, the methods evaluated within a single paper are often limited in scope, e.g., newly proposed approaches from the same period are rarely compared empirically. Secondly, although traffic prediction models can also be adapted for traffic data imputation, systematic comparisons between imputation and prediction models under consistent settings are scarce. Thirdly, most evaluations only report overall average errors, neglecting performance analysis during difficult periods with significant traffic fluctuations, which obscures differences in model performance under such scenarios. Thirdly, model efficiency analysis, a key factor for real-world deployment, is frequently neglected.
To meet the need, we design an unified benchmarking pipeline to evaluate ten representative or recently proposed imputation models for spatial-temporal traffic data. Our evaluation spans both traffic flow and speed datasets, covering all possible missing patterns across spatial and temporal dimensions, accompanied by detailed performance analysis. Meanwhile, to provide readers with an overall picture of the latest advancements on traffic data imputation and to enhance their insights into model characteristics, we offer a holistic overview of imputation models for traffic data and categorize various methods from a practice-oriented perspective.
@@
While there are several recent surveys on data imputation models, they pay little attention to their applications in spatial-temporal traffic scenarios. For instance, Miao et al. [1] provides a general discussion on imputation across various data types (including numerical, categorical, and mixed data). Fang et al. [2] focuses on introducing deep learning-based imputation models for time series without empirical evaluations. Wang et al. and Du et al. review deep learning methods for multivariate time series imputation and introduce benchmark platforms [3, 4], yet they exclude imputation approaches beyond deep neural networks, such as tensor completion methods. Compared to them, this paper distinguishes itself in two key aspects. Firstly, we emphasize representative and emerging imputation models specifically applicable for spatial-temporal traffic data. Secondly, beyond simply analyzing overall experimental results, we provide visualizations and delve into a detailed performance analysis during critical difficult periods, addressing core concerns within the ITS domain. The contributions of this work are summarized as follows:
@@
    Comprehensive evaluation. We conduct extensive experiments on 10 representative models across three traffic datasets, evaluating their performance under 20 scenarios (comprising combinations of 4 missing patterns and 5 missing rates). A detailed analysis of both effectiveness and efficiency is provided to serve as a practical guideline for model selection and application.
The rest of this paper is organized as follows: Section II reviews related work. Section III gives an introduction of the spatial-temporal traffic data imputation problem. Section IV presents a overview on imputation models and proposes the taxonomy. Section V introduces the evaluated models. Section VI presents the unified benchmarking pipeline and experimental results. Section VII concludes the paper.
II Related Work
In this section, we review four categories of works relevant to spatial-temporal traffic data imputation. Specifically, multi-variable time series imputation is a broad topic, with spatial-temporal graph data imputation as its subbranch. Spatial-temporal graph imputation models are the most suitable kind of approaches to achieve traffic data imputation. Besides, spatial-temporal kriging is a special case of spatial-temporal graph imputation when there are continuous missing data in the temporal dimension. Time series probabilistic imputation further incorporates uncertainty for the imputation values.
II-A Muti-variable Time Series Imputation
II-B Spatial-temporal Graph Imputation
In contrast to MTSI, the spatial-temporal graph imputation (STGI) approach [46, 47, 48, 49, 50] incorporates an extra graph structure to explicitly represent the spatial correlations inherent in time series data. Typically, STGI utilizes recurrent neural networks (RNNs) and attention mechanisms to capture temporal dependencies, combined with graph neural networks (GNNs) to capture spatial dependencies. For example, GRIN [51] utilizes a specialized message-passing GNN to effectively capture spatial dependencies and impute time series in both forward and backward directions. SPIN [52] incorporates sparse spatial-temporal attention mechanisms, which only propagate observed data points, thus efficiently diminishing noise associated with missing values. MDGCN [53] strengthens the modeling of spatial correlations by learning the graph structure and integrates an external memory network to retain global spatiotemporal knowledge. GCASTN [54] introduces a generative-contrastive self-supervised learning framework, enhancing model robustness by comparing two different augmented perspectives. ImputeFormer [55] presents a Transformer-based architecture that leverages low-rankness principles to bridge the gap between deep learning and tensor completion methods for data imputation.
II-C Spatial-temporal Kriging
II-D Time Series Probabilistic Imputation
The imputation above methods are focused on deterministic imputation, where only a plausible value is filled in for the missing data. In contrast, Time Series Probabilistic Imputation (TSPI) seeks to provide a range for the missing values, thereby quantifying the confidence level in the imputation [35]. A typical approach in this domain involves using generative models to capture the data distribution. For instance, GP-VAE [69] utilizes a Gaussian process as a prior to represent temporal dependencies.
III Preliminaries
@@ III-A Problem Statement
[…]
@@
$\delta_{v,t}=\begin{cases}0&\text{ if }t=1\\ 1&\text{ if }t>1\text{ and }m_{v,t-1}=1\\ \delta_{v,t-1}+1&\text{ if }t>1\text{ and }m_{v,t-1}=0\end{cases}$ (1)
@@ III-B Missing Patterns
Figure 1: Illustrations of four missing patterns in traffic data, where white grid indicates missing data and gray grid indicates observed data. Specifically, SRTR may be caused by random signal or network interruption, SCTR may be caused by equipment failure over a group of devices due to some factors like network interruption in an area, SRTC may be caused by some sensors experiencing equipment failure, or network outages over a period of time, and SCTC may be caused by some reasons like power failure on a group of devices over a period of time in an area.
IV Models Summary
In this section, we start by providing a concise overview on models for time series imputation, within which traffic data imputation is a special case. Next, we introduce a practice-oriented taxonomy for traffic data imputation models, emphasizing their commonalities.
IV-A Overview on Imputation Models for Time series
To provide a overall picture on research popularity for time series imputation models, we present statistics on the annual number of related published papers, as shown in Figure 2. We also summarize the imputation model on key techniques, training methods, tasks, and dataset as shown in Table I. Specifically, we search for imputation-related papers in the IEEE Xplore, ACM digital libraries, and Web of Science. Our statistics consider the papers published from 2018 to 2024 in prestigious venues, e.g., renowned conferences and journals, including NeurIPS, ICLR, AAAI, ICML, IJCAI, ICDE, VLDB, KDD, TKDE, TITS. To ensure the papers focus on time series imputation, we use keywords including“missing data,” “imputation data,” and “interpolation” during our search. Finally, based on these papers, we conduct a manual filtering process to select the papers on time series imputation.
Figure 2: The annual number of papers published on time series imputation from 2018 to 2024.
@@ IV-A Overview on Imputation Models for Time series
Reference Model Tasks Key Techniques Training Method Dataset Cites
@@
$\bigstar$ Cao et al. - NeurIPS’18 [74] BRITS MTS imputation RNN Self-supervised Air Quality/ Health-care/ Human Activity 426
Shao et al. - TITS’18 [11] CTD STG imputation MF self-supervised Traffic 18
Kim et al. - IJCAI’18 [15] TBF MTS imputation RNN self-supervised Health-care 65
@@
Luo et al. - NeurIPS’18 [18] GRUI-GAN MTS imputation RNN GAN Air Quality/ Health-care 402
Yoon et al. - ICML’18 [35] GAIN Probabilistic imputation GAN GAN UCI 859
@@
Che et al. - ICML’18 [19] MR-HDMM MTS imputation RNN self-supervised Health-care/Air Quality 43
@@
Yoon et al. - ICLR’18 [20] M-RNN MTS imputation RNN self-supervised Health-care 30
@@
Wang et al. - TITS’19 [56] OCC ST Kriging RNN self-supervised Traffic 18
Li et al. - TITS’19 [48] MVLM STG imputation RNN self-supervised Traffic 151
Wang et al. - TITS’19 [12] TAS-LR STG imputation MF unsupervised Traffic 73
@@
$\bigstar$ Luo et al. - IJCAI’19 [34] E2GAN Probabilistic imputation RNN GAN Air Quality/Health-care 166
Liu et al. - NeurIPS’19 [43] NAOMI Probabilistic imputation RNN GAN Traffic 102
Hwang et al. - ICML’19 [37] Hexagan Probabilistic imputation GAN GAN/semi-supervised Images/Synthetic 44
Li et al. - ICLR’19 [36] Misgan Probabilistic imputation GAN GAN Images 192
@@
Ma et al. - TKDE’19 [21] OSICM MTS imputation kNN self-superviesd Air Quality/Health-care/Wiki4HE 13
@@
Zhang et al. - ICDE’19 [22] IIM MTS imputation KNN self-supervised UCI/KEEL/Siemens 37
@@
Wu et al. - ICDE’19 [23] HKMF-T MTS imputation MF self-supervised Traffic 8
@@
Park et al. - KDD’19 [38] Imp-GAIN MTS imputation RNN+Attention GAN Health-care 12
Jiang et al. - ICML’20 [24] BiLSTM-A MTS imputation RNN self-supervised Air Quality 1
@@
Khayati et al. - VLDB’20 [7] ORBITS MTS imputation Centroid Decomposition self-supervised Soccer/MotionSense/BAFU/Gas 5
Ma et al. - KDD’20 [25] MIDIA MTS imputation AutoEncoder self-supervised Air Quality/Adult/Car 17
@@
$\bigstar$ Shukla et al. - ICLR’21 [44] mTAN MTS imputation Attention supervised Clinical /Human Activity 86
@@
Gong et al. - IJCAI’21 [13] SMV-NMF STG imputation k-means + MF unsupervised Urban Statistical 20
@@
Xiao et al. -TITS’21 [39] MTCIU GAN MTS imputation GAN GAN Traffic 3
@@
$\bigstar$ Chen et al. - TITS’21 [14] LATC MTS imputation TC unsupervsed Traffic 47
@@
Zhang et al. - TITS’21 [40] SA-GAIN MTS imputation Attention GAN Traffic 27
@@
Deng et al. - TITS’21 [49] GTC STG imputation TC unsupervsed Traffic 16
Xu et al. - TITS’21 [26] GA-GAN Probabilistic imputation GraphSAGE GAN Traffic 5
@@
Miao et al. - AAAI’21 [27] SSGAN MTS imputation RNN semi-supervised Human activity/Clinic/Meteorologic 54
@@
Tashiro et al. - NIPS’21 [70] CSDI Probabilistic imputation Attention + Diffusion Model self-supervised Air Quality / Health 118
Dai et al. - ICML’21 [41] MI-GAN Probabilistic imputation GAN GAN ADNI/Synthetic 11
Qin et al. - KDD’21 [46] ST-SCL STG imputation VAE semi-supervised Traffic 16
@@
Liang W et al. - TITS’22 [57] STAR ST Kriging GNN+Attention supervised Traffic 50
@@
Lei M et al. - TITS’22 [58] BKMF ST Kriging GP unsupervised Health-care 11
@@
Wang A et al. - TITS’22 [47] GSTAE STG Imputation GCN+GRU supervised Traffic 0
Li J et al. - TKDE’22 [17] MT-CSR MTS Imputation CNN supervised Traffic 2
@@
Luo X et al. - TKDE’22 [28] NeuLFT MTS Imputation TC unsupervised Dynamic interaction networks 63
Zhao Y et al. - AAAI’22 [29] GC MTS Imputation EM unsupervised Synthetic 6
@@
Fan J. - AAAI’22 [10] D-NLMC MTS Imputation TC unsupervised Synthetic Air Quality / Temperature / Chlorine Level Dataset 1
@@
Marisca I et al. - NIPS’22 [52] SPIN MTS Imputation Attention supervised Traffic 9
@@
Ipsen N B et al. - ICLR’22 [75] supMIWAE Probabilistic Imputation Autoencoder supervised IMAGE / CLASSIFICATION 29
@@
Liu Y et al. - ICLR’22 [30] MBMF Probabilitisc Imputation MissForest supervised Synthetic / Medical 3
Lao D et al. - KDD’22 [59] WSGNN ST Kriging GCN weakly-supervised Cora / Citeseer / Pubmed 8
Xing J et al. - TITS’23 [8] DFCP MTS imputation TC unsupervised CL / LPR 0
@@
Xiuqin Xu et al. - TITS’23 [9] HRST-LR MTS imputation TC unsupervised Traffic 0
Varga B et al. - TITS’23 [60] STK ST Kriging Kriging unsupervised Traffic 1
@@
Blázquez-García A et al. - TKDE’23 [31] MGP MTS imputation GP unsupervised Synthetic / CLASSIFICATION 0
@@
Chowdhury R R et al. - AAAI’23 [45] PrimeNet MTS imputation Attention self-supervised PhysioNet / MIMIC-III 0
@@
Zhang X et al. - ICML’23 [16] UTDE-mTAND MTS imputation Attention supervised MIMIC-III 1
Zhao H et al. - ICML’23 [32] TDM MTS imputation NN unsupervised UCI 10

### v2

# A Survey and Benchmarking of Spatial-Temporal Traffic Data Imputation Models

@@ Abstract
Traffic data imputation is a critical preprocessing step in intelligent transportation systems, underpinning the reliability of downstream transportation services. Despite substantial progress in imputation models, model selection and development for practical applications remains challenging due to three key gaps: 1) the absence of a model taxonomy for traffic data imputation to trace the technological development and highlight their distinct features. 2) the lack of unified benchmarking pipeline for fair and reproducible model evaluation across standardized traffic datasets. 3) insufficient in-depth analysis that jointly compare models across multiple dimensions, including effectiveness, computational efficiency and robustness. To this end, this paper proposes practice-oriented taxonomies for traffic data missing patterns and imputation models, systematically cataloging real-world traffic data loss scenarios and analyzing the characteristics of existing models. We further introduce a unified benchmarking pipeline to comprehensively evaluate 11 representative models across various missing patterns and rates, assessing overall performance, performance under challenging scenarios, computational efficiency, and providing visualizations. This work aims to provide a holistic perspective on traffic data imputation and to serve as a practical guideline for model selection and application in intelligent transportation systems.
@@ I Introduction
Traffic data is an example of spatial-temporal graph sequences, in which each node represents a traffic sensor that continuously collects time-series observations, and the edges represent the relationships between sensors. Fig. 1 illustrates a real-world highway scenario where traffic data are collected through various sensors, such as loop detectors, radar detectors, and video detectors. These data exhibit strong spatial and temporal correlations. For example, traffic data from three connected sensors share similar changing trends, while a single traffic data sequence shows temporal autocorrelation and periodicity.
Figure 1: An illustration showing how traffic data can be modeled as spatial–temporal graph sequences.
@@
As a fundamental resource for Intelligent Transportation Systems (ITS) [1], it underpins key downstream services, such as traffic management [2], traffic data prediction [3, 4, 5], and traffic signal control [6, 7, 8]. Despite the increasing deployment of sensors to collect real-time traffic information, missing data remains a persistent issue due to various factors like equipment malfunctions, power outages, and network failures. These missing data hinder ITS to accurately perceive traffic changes over time, resulting in the performance decline of downstream transportation services. Therefore, traffic data imputation is a crucial preprocessing step in ITS, ensuring the quality of traffic data.
@@
Numerous models have been specifically developed to solve this problem, focusing on two main challenges: 1) capturing the dynamics of traffic data with missing values across both spatial and temporal dimensions. 2) training models without access to ground-truth observations for the missing entries. In addition to traffic-specific imputation methods, general multivariate time series imputation models can also be adapted to this task. Despite significant progress in this field, developing an effective model for specific real-world scenarios in ITS remains difficult due to three key factors.
    Absence of a comprehensive survey and practical taxonomy. This gap makes it difficult to trace the technological development of traffic data imputation models and to effectively apply emerging techniques to this task. A thorough survey is needed to summarize existing models and highlight their characteristics and strengths. Moreover, a taxonomy is required to clarify the rationale behind different designs and to guide the development of new models. In addition, studies on traffic forecasting [5, 9, 10], which are closely related to imputation and share transferable techniques, should also be incorporated and discussed.
    Lack of a unified benchmarking pipeline. Existing studies often evaluate models on different datasets and only a limited set of missing patterns. These missing patterns describe how traffic data are lost, such as randomly or continuously along spatial and temporal dimensions. However, few empirical studies compare models on standardized spatiotemporal traffic datasets with comprehensive missing patterns, making fair comparisons difficult. Additionally, although both traffic flow and traffic speed are critical for ITS, few studies have conducted experiments on both simultaneously.
@@
    Limited in-depth evaluation and analysis. First, methods evaluated within a single paper are often limited in scope; for example, newly proposed approaches from the same period or based on different technical architecture are rarely compared empirically. Second, most evaluations report only overall average errors, neglecting performance during periods with significant traffic fluctuations, which obscures model differences under challenging scenarios. Third, model efficiency, an important factor for real-world deployment, is frequently overlooked.
To address these challenges and advance research and applications in traffic data imputation, we first present a holistic overview of existing models and categorize them from a practice-oriented perspective. This provides readers with a comprehensive understanding of recent advancements and deeper insights into model characteristics. Furthermore, we design a unified benchmarking pipeline to evaluate 11 representative models for spatio-temporal traffic data imputation, each regarded as classical within its respective technical paradigm. Our evaluation covers both traffic flow and speed datasets, encompasses diverse missing patterns across spatial and temporal dimensions, and includes detailed performance analyses. Through this, we aim to guide readers in selecting and deploying suitable models for practical scenarios.
@@
While there are several recent surveys on data imputation models, they pay little attention to their applications in spatial-temporal traffic scenarios. For instance, Miao et al. [11] provides a general discussion on imputation across various data types (including numerical, categorical, and mixed data). Fang et al. [12] focuses on introducing deep learning-based imputation models for time series without empirical evaluations. Wang et al. [13] and Du et al [14] review deep learning methods for multivariate time series imputation and introduce benchmark platforms, yet they exclude imputation approaches beyond deep neural networks, such as tensor completion methods. Zhang et al. [15] review public spatial-temporal traffic datasets and commonly used approaches, categorizing imputation models into interpolation-based, statistical learning-based, and prediction-based methods. However, their taxonomy is overly coarse, making it difficult to capture the commonalities and distinctive characteristics of different models. In addition, they do not provide experimental evaluations to assess which technologies most effective for different missing-data scenarios. Compared to them, our paper distinguishes itself in two key aspects. Firstly, we focus on the models imputation specifically applicable for spatial-temporal traffic data imputation. Secondly, beyond merely surveying these models, we conduct experiments to compare representative approaches. Rather than relying solely on overall performance metrics, we provide a detailed analysis of model behavior during critical and challenging periods, directly addressing practical concerns in the ITS domain.
The contributions of this work are summarized as follows:
@@
    Comprehensive evaluation and in-depth analysis. We conduct extensive experiments on 11 representative models across four traffic datasets, evaluating their performance under 20 scenarios (comprising combinations of 4 missing patterns and 5 missing rates). A detailed analysis of both effectiveness and efficiency is provided to serve as a practical guideline for model selection and application.
The rest of this paper is organized as follows: Section II gives an introduction of the spatial-temporal traffic data imputation problem. Section III provides an overview of existing time-series imputation models and reviews the literature to position traffic data imputation within the broader context of related tasks. It then proposes a practice-oriented taxonomy to clarify the relationships and distinctions among these models. Section IV introduces the imputation models evaluated in this paper. Section V presents the unified benchmarking pipeline and experimental results. Section VI concludes the paper.
II Preliminaries
@@ II-A Problem Statement
Traffic Network. We define the traffic network as a graph $G=(V,E,A)$, where $V$ represents the set of $|V|=N$ nodes (e.g., loop detectors or video cameras deployed on the traffic network). $E$ represents the set of edges that connect nodes together. $A\in{\mathbb{R}^{N\times N}}$ is the adjacency matrix representing the proximity between nodes.
@@
Figure 2: Illustrations of four missing patterns in traffic data, where white grids indicate missing data and gray grids indicate observed data. Specifically, SRTR may be caused by random signal or network interruption, SCTR may be caused by equipment failure over a group of devices due to some factors like network interruption in an area, SRTC may be caused by some sensors experiencing equipment failure, or network outages over a period of time, and SCTC may be caused by some reasons like power failure on a group of devices over a period of time in an area.
@@
[…]
@@ II-B Missing Patterns
III Model Taxonomy
In this section, we start by providing a concise overview on models for time series imputation, within which traffic data imputation is a special case. Next, we introduce a practice-oriented taxonomy for traffic data imputation models to elucidate the relationships and distinctions among them.
III-A Overview on Imputation Models for Time series
To provide a overall picture on research popularity for time series imputation models, we present statistics on the annual number of related published papers, as shown in Figure 3. Specifically, we search for imputation-related papers in the IEEE Xplore, ACM digital libraries, and Web of Science. Our statistics consider the papers published from 2018 to 2025 in prestigious venues, e.g., renowned conferences and journals, including ICDE, VLDB, SIGMOD, KDD, TKDE, NeurIPS, ICLR, AAAI, ICML, IJCAI, TITS. To ensure the papers focus on time series imputation, we use keywords including “missing data,” “imputation data,” and “interpolation” during our search. Finally, based on these papers, we conduct a manual filtering process to select the papers on time series imputation.
Generally, studies on time-series imputation can be grouped into four main categories. In particular, multivariate time-series imputation is a broad area in which spatial–temporal graph data imputation is often treated as a subfield. Spatial–temporal graph imputation models represent the most suitable class of approaches for traffic data imputation. Moreover, spatial–temporal kriging can be viewed as a special case of spatial–temporal graph imputation, particularly for scenarios with continuous temporal gaps. Finally, probabilistic time-series imputation methods further estimate the uncertainty associated with the imputed values. A brief introduction to each of these four categories is given below.
Figure 3: The annual number of papers on time series imputation published from 2018 to 2025.
III-A1 Muti-variable Time Series Imputation
III-A2 Spatial-Temporal Graph Imputation
In contrast to MTSI, spatial-temporal graph imputation (STGI) [57, 58, 59, 60, 61] incorporates an extra graph structure to explicitly represent the spatial correlations inherent in time series data. Typically, STGI utilizes recurrent neural networks (RNNs) and attention mechanisms to capture temporal dependencies, combined with graph neural networks (GNNs) to capture spatial dependencies. For example, GRIN [62] utilizes a specialized message-passing GNN to effectively capture spatial dependencies and impute time series in both forward and backward directions. SPIN [63] incorporates sparse spatial-temporal attention mechanisms, which only propagate observed data points, thus efficiently diminishing noise associated with missing values. MDGCN [64] strengthens the modeling of spatial correlations by learning the graph structure and integrates an external memory network to retain global spatiotemporal knowledge. GCASTN [65] introduces a generative-contrastive self-supervised learning framework, enhancing model robustness by comparing two different augmented perspectives. ImputeFormer [66] presents a Transformer-based architecture that leverages low-rankness principles to bridge the gap between deep learning and tensor completion methods for data imputation.
III-A3 Spatial-Temporal Kriging
III-A4 Time Series Probabilistic Imputation
Uncertainty quantification and probabilistic modeling for traffic data  [80, 81] is vital in practice. Most imputation methods mentioned above are focused on deterministic imputation, where only a plausible value is filled in for the missing data. In contrast, Time Series Probabilistic Imputation (TSPI) seeks to provide a range for the missing values, thereby quantifying the confidence level in the imputation [46]. A typical approach in this domain involves using generative models to capture the data distribution. For instance, GP-VAE [82] utilizes a Gaussian process as a prior to represent temporal dependencies.
@@ III-A4 Time Series Probabilistic Imputation
Reference Model Tasks ST Model Loss Designs Dataset
Techniques Model Types Training Strategies
@@
$\bigstar$ Cao et al. - NeurIPS’18 [87] BRITS MTS imputation RNN Predictive (MAE) Reconstruction-based SSL Air Quality/ Health-care/ Human Activity
Shao et al. - TITS’18 [22] CTD STG imputation TC Predictive Reconstruction-based SSL Traffic
Kim et al. - IJCAI’18 [26] TBM MTS imputation RNN Predictive Reconstruction-based SSL Health-care
@@
Luo et al. - NeurIPS’18 [29] GRUI-GAN MTS imputation RNN Generative (GAN) Masked SSL Air Quality/ Health-care
Yoon et al. - ICML’18 [46] GAIN Probabilistic imputation RNN Generative (GAN) Masked SSL UCI
@@
Che et al. - ICML’18 [30] MR-HDMM MTS imputation RNN Generative (VAE) Reconstruction-based SSL Health-care/Air Quality
@@
Yoon et al. - ICLR’18 [31] M-RNN MTS imputation RNN Predictive (RMSE) Masked SSL Health-care
@@
Wang et al. - TITS’19 [67] OCC ST Kriging RNN Predictive (RMSE) Reconstruction-based SSL Traffic
Li et al. - TITS’19 [59] MVLM STG imputation RNN Predictive (MAE) Reconstruction-based SSL Traffic
Wang et al. - TITS’19 [23] TAS-LR STG imputation TC Predictive Masked SSL Traffic
@@
$\bigstar$ Luo et al. - IJCAI’19 [45] E2GAN Probabilistic imputation RNN Generative (GAN) Masked SSL Air Quality/Health-care
Liu et al. - NeurIPS’19 [54] NAOMI Probabilistic imputation RNN Generative (GAN) Masked SSL Traffic
Li et al. - ICLR’19 [47] Misgan Probabilistic imputation RNN Generative (GAN) Masked SSL Images
@@
Ma et al. - TKDE’19 [32] OSICM MTS imputation KNN Predictive (RMSE) Reconstruction-based SSL Air Quality/Health-care/Wiki4HE
@@
Zhang et al. - ICDE’19 [33] IIM MTS imputation KNN Predictive (RMSE) Reconstruction-based SSL UCI/KEEL/Siemens
@@
Wu et al. - ICDE’19 [34] HKMF-T MTS imputation TC Predictive (RMSE) Reconstruction-based SSL Traffic
@@
Park et al. - KDD’19 [49] Imp-GAIN MTS imputation RNN+Attention Generative (GAN) Reconstruction-based SSL Health-care
Jiang et al. - ICML’20 [35] BiLSTM-A MTS imputation RNN Predictive Masked SSL Air Quality
@@
Khayati et al. - VLDB’20 [18] ORBITS MTS imputation TC Predictive (RMSE) Reconstruction-based SSL Soccer/MotionSense/BAFU/Gas
@@
$\bigstar$ Shukla et al. - ICLR’21 [55] mTAN MTS imputation RNN + Attention Generative (VAE) Reconstruction-based SSL Clinical /Human Activity
@@
Gong et al. - IJCAI’21 [24] SMV-NMF STG imputation TC Predictive (RMSE) Reconstruction-based SSL Urban Statistical
@@
Xiao et al. -TITS’21 [50] MTCIU GAN MTS imputation RNN Generative (GAN) Reconstruction-based SSL Traffic
@@
$\bigstar$ Chen et al. - TITS’21 [25] LATC MTS imputation TC Predictive (RMSE) Masked SSL Traffic
@@
Zhang et al. - TITS’21 [51] SA-GAIN MTS imputation Attention Generative (GAN) Masked SSL Traffic
@@
Deng et al. - TITS’21 [60] GTC STG imputation TC Predictive (RMSE) Reconstruction-based SSL Traffic
Xu et al. - TITS’21 [37] GA-GAN Probabilistic imputation GNN Generative (GAN) Reconstruction-based SSL Traffic
$\bigstar$ Wu et al. - AAAI’21 [73] IGNNK STG imputation GNN Predictive (RMSE) Reconstruction-based SSL Traffic / Solar power
@@
Miao et al. - AAAI’21 [38] SSGAN MTS imputation RNN Generative (GAN) Semi-Supervised Learning Human activity/Clinic/Meteorologic
@@
Tashiro et al. - NIPS’21 [83] CSDI Probabilistic imputation Attention Generative (Diffusion) Masked SSL Air Quality / Health
Dai et al. - ICML’21 [52] MI-GAN Probabilistic imputation NN Generative (GAN) Masked SSL ADNI/Synthetic
Qin et al. - KDD’21 [57] ST-SCL STG imputation GNN Generative (VAE) Reconstruction-based SSL Traffic
@@
Liang et al. - TITS’22 [68] STAR ST Kriging GNN + Attention Predictive (RMSE) Reconstruction-based SSL Traffic
@@
Lei et al. - TITS’22 [69] BKMF ST Kriging TC Predictive (RMSE) Reconstruction-based SSL Health-care
@@
Wang et al. - TITS’22 [58] GSTAE STG Imputation GNN + RNN Predictive (RMSE) Masked SSL Traffic
Li et al. - TKDE’22 [28] MT-CSR MTS Imputation CNN Predictive (RMSE) Reconstruction-based SSL Traffic
@@
Luo et al. - TKDE’22 [39] NeuLFT MTS Imputation TC Predictive (RMSE) Reconstruction-based SSL Dynamic interaction networks
@@
Fan et al. - AAAI’22 [21] D-NLMC MTS Imputation TC Predictive (RMSE) Reconstruction-based SSL Synthetic Air Quality / Temperature / Chlorine Level Dataset
@@
Marisca et al. - NIPS’22 [63] SPIN MTS Imputation GNN + Attention Predictive (MAE) Masked SSL Traffic
@@
Ipsen et al. - ICLR’22 [88] supMIWAE Probabilistic Imputation NN Predictive Supervised Learning IMAGE / CLASSIFICATION
@@
Liu et al. - ICLR’22 [41] MBMF Probabilitisc Imputation Causal Discovery Predictive (RMSE) Reconstruction-based SSL Synthetic / Medical
Xing et al. - TITS’23 [19] DFCP MTS imputation TC Reconstruction-based SSL Reconstruction-based SSL CL / LPR
@@
Xu et al. - TITS’23 [20] HRST-LR MTS imputation TC Predictive (RMSE) Reconstruction-based SSL Traffic
@@
Blázquez-García et al. - TKDE’23 [42] MGP MTS imputation RNN Predictive (RMSE) Reconstruction-based SSL Synthetic / CLASSIFICATION
@@
Chowdhury et al. - AAAI’23 [56] PrimeNet MTS imputation Attention Predictive Masked SSL PhysioNet / MIMIC-III
@@
Zhang et al. - ICML’23 [27] UTDE-mTAND MTS imputation Attention Predictive Reconstruction-based SSL MIMIC-III